OpenStudy (anonymous):

A family's income should be approximetly 2/5 of the price of a house. If your family's income is $36,000, what is the cost of a house you could afford to buy?

OpenStudy (kittiwitti1):

sorry wrote that wrong - deleted it

OpenStudy (ribhu):

couldn't send from here so dropped in ur inbox

OpenStudy (ribhu):

\[\let the price of the house be x so income becomes 2/5(x) but income is given as : $36000 so 2/5 (x) = 36000 x = 36000*5/2 x= 90000 hence price of the house is $90000\]

OpenStudy (kittiwitti1):

Ok. So income is 2/5 (or 20%, 0.2) of house price. We have the income... which is 2/5 of unknown house price. If we had the house price and not the income, we would take 0.2*[house price]. But we have the income. So let's reverse that process.

OpenStudy (kittiwitti1):

Sorry - not 20%, 40% It's 0.4

OpenStudy (kittiwitti1):

Equation:\[\frac{\text{income value }x}{0.4}\]

OpenStudy (kittiwitti1):

Plug it in, what did you get @katlin95 ? :)

OpenStudy (anonymous):

90,000 for my answer!

OpenStudy (kittiwitti1):

Yes :)

OpenStudy (kittiwitti1):

I have to sleep now, but good luck on the rest of your work~ ^^ Night! (:

OpenStudy (anonymous):

thanks!!

OpenStudy (kittiwitti1):

No problem :D